# Scalable modular architecture for universal quantum computation

### v2

@@ I Introduction
The development of quantum hardware is making fast progress over the past few years, bringing us from the NISQ era [1] closer to fault tolerance [2] and practical quantum error correction [3, 4]. At the forefront are architectures based on trapped neutral [4, 5] and ionic [2, 6, 7] atoms as well as superconducting qubits [3, 8, 9, 10, 11, 12], all of which are expected to scale up further. Important landmarks such as the implementation of fault-tolerant operations [2] or demonstrations of break-even for quantum error correction [3, 4] and quantum simulation [13, 7] shift the focus to larger devices. Typically, scalable designs are obtained via a modular approach, linking together multiple smaller quantum processors with limited connectivity rather than developing larger processors [14, 15, 16, 17].
@@
One requirement for realizing quantum computing is the ability to perform every unitary operation [18]. In control theory, this is termed evolution operator controllability [19]. Controllability analysis can thus be used as a valuable tool in quantum chip design to ensure that all operations on the device are possible while at the same time simplifying the device [20, 21]. The same concepts that are at the core of controllability analysis can also be employed to the expressivity of parameterized quantum circuits [21, 22]. Controllability tests can be used to identify the sweet spot balancing universality and the complexity of the system’s architecture [20] where complexity refers to qubit-qubit couplings and external qubit controls, i.e., the resources to connect and alter the qubits. Reducing the amount of built-in resources while maintaining the computational capability of the quantum device has not yet been a focus for device design but will be important for further scaling up devices, as is evident just by looking at the number of classical control lines needed to operate today’s most advanced QPUs [2, 3, 4, 5, 6, 8, 9, 10].
@@
Controllability can be rigorously proven by studying the dynamical Lie algebra of the system, i.e., the algebra generated by a system’s Hamiltonian split into its time-independent drift and all the control operators [19]. Full rank of the dynamical Lie algebra implies that the system is evolution operator controllable [19]. However, only in exceptional cases can the dynamical Lie algebra be calculated analytically [23, 24, 25, 26, 27, 28]. Alternatively, one can determine the rank of the dynamical Lie algebra numerically. The most straightforward way would be to construct a complete basis of the dynamical Lie algebra (which is also a vector space) but this becomes unstable in the required orthogonalization step already for a relatively small number of qubits. Construction of the full algebra can be circumvented by graph methods [29, 20]. Nevertheless, this approach will eventually fail too, due to the exponential scaling, with the number of qubits, of the size of Hilbert space and thus the dimension of the graph. While the curse of dimensionality can — in principle — be evaded by carrying out the controllability test via a parametric quantum circuit in a hybrid approach combining classical optimization with a quantum device [21], it remains yet to be seen how reliably the dimensional expressivity test at the core of this method can be carried out on noisy quantum devices and for larger numbers of qubits. Thus, at present, controllability tests that are generally applicable as well as demonstratedly reliable and scalable are missing. A change in perspective is made possible when proving controllability of a large quantum system from the controllability of its constituent subsystems [30, 31, 28] .
@@
Here, we leverage this change in perspective for quantum chip design. Instead of attempting to prove controllability for all qubits at once, we show how to build up controllable qubit arrays from smaller, fully controllable subsystems. We can start from building blocks as small as two (or a few) qubits where it is straightforward to prove controllability. Surprisingly, a single controllable entangling link between two qubits in each of the quantum processing subunits is sufficient. Such links are routinely realized in current quantum computing architectures, for example via tunable couplers [32, 33, 34] for superconducting qubits. While, for a given number of qubits, our approach will not result in the QPU with the smallest possible number of local controls and qubit-qubit couplings, it allows for designs with significantly reduced resources. As an example, we first discuss how two five-qubit arrays that are connected via a tunable coupler result in a controllable ten-qubit array. We then show that this method can be scaled up to generate large devices, thus opening a new route for more resource-efficient device design.
@@ II Controllability of bipartite systems with an entangling control
We briefly recap evolution operator controllability of a quantum system [19]. Consider a system with Hilbert space dimension $\dim(\mathcal{H})=n$ whose Hamiltonian can be written as
@@
Now, we show how to ensure that controllability is maintained when assembling a larger system by connecting smaller controllable systems. The main idea is rather simple. Assume that we have two, possibly different, qubit arrays (”modules”) with given configurations of qubit-qubit couplings and local qubit controls, that are both evolution operator controllable. Then it suffices to connect the two arrays with a control acting on two qubits, one in each module, that is capable of generating entanglement between them, for the resulting multipartite system (composed of the original arrays and the new entangling control) to also be evolution operator controllable. We can keep connecting modules to the original system with one entangling operation per new module and thus design an arbitrarily large system that is evolution operator controllable.
@@
This result is a particular case of more general results on controllability of multipartite systems [30, 31]. Our proof of Theorem 1, presented in Appendix A, builds a full basis of the Lie algebra space. This new constructive method provides further information about the depth of commutators required to implement certain elements in the Lie algebra, which can in principle be related to how easily some unitary operations can be implemented on the global system.
@@
Considering the simplest example from a gate-oriented perspective, we recover a well-known statement as special case of Theorem 1 — for two qubits, the set of local operations together with any entangling two-qubit gate form a universal set. Here, having access to all local gates implies that the two original systems, the separate qubits, are evolution operator controllable. The entangling gate has the same effect as the entangling control that we use to connect the systems. Note that already in the bipartite case, Theorem 1 is slightly more general: Substitute qudits for the qubits, perhaps of different dimensions $d_{1}=2^{M}$ and $d_{2}=2^{N}$. Since the study is restricted to qubit arrays, the dimension of the qudits is also restricted to powers of 2. The set of local unitary operators for two qubits $SU(2)\otimes SU(2)$ is replaced by $SU(d_{1})\otimes SU(d_{2})$. Every local gate in $SU(d_{1})\otimes SU(d_{2})$ can be implemented given the assumption of controllability on the separate partitions. Theorem 1 then states that given two controllable qudits with dimensions $d_{1}$ and $d_{2}$ and an entangling Hermitian operator $\hat{G}$, the set composed of local operations on the two qudits and the entangling rotation gates $\hat{R}_{\hat{G}}(\phi)$ form a universal set. This universality of the set of local gates plus an entangling gate on qudits of the same size is a lesser known fact in quantum information [35, 36, 37].
@@
Once two controllable systems are connected by an entangling control (e.g. a tunable coupler), they can be understood as a controllable system in itself. This implies that we can connect it to a different controllable system using a new entangling control to build a larger yet controllable system. Note that the type of entangling operation between partitions is not relevant to prove controllability. In particular, the statement is true for any two-qubit tunable coupling as long as each of the qubits belongs to a different partition.
@@
Naturally, having only one entangling operation between two partitions will limit the speed at which information (and in particular entanglement) can travel through the system. At the same time, a modular approach will also ease transpilation of quantum algorithms, making the use of partitions that are large enough to run certain subtasks a viable option. This type of architecture is also ideal for using each partition as sub-unit when parallelizing a more complex quantum algorithm. While these use cases would be specifically tailored to the modular design, the universality of the composite system ensures its use it for any kind of quantum circuit.
@@
Figure 1: (a): Controllable five-qubit system using just two controls [20]. Qubits with (without) local controls are shaded (blank) and static couplings are drawn as straight black lines. (b): Controllable system built up from two copies of the system in (a) with a tunable coupling (blue zig-zag connection).
@@
As an example, consider a T-shaped five-qubit system, inspired by the former five-qubit ibm_quito quantum processor by IBM (cf. https://quantum.cloud.ibm.com/docs/en/guides/retired-qpus), the controllability of which has been proven using a graph method [20]. When properly placed, two local controls are sufficient for the five-qubit system with fixed two-qubit couplings in the form of $\hat{H}^{(k,l)}=\hat{X}^{(k)}\hat{X}^{(l)}+\hat{Y}^{(k)}\hat{Y}^{(l)}$ to be evolution operator controllable, as illustrated also in Fig. 1(a). Its Hamiltonian is given by
$\hat{H}^{5Q}(t)=$ $-\sum_{j=0}^{4}\frac{\omega_{j}}{2}\hat{Z}^{(j)}+\sum_{\begin{subarray}{c}(k,l)\in\{(0,1),(1,2),\\ (1,3),(3,4)\}\end{subarray}}J_{k,l}\hat{H}^{(k,l)}$
$+u_{1}(t)\hat{X}^{(1)}+u_{2}(t)\hat{X}^{(3)}\,$, (3)
@@
Connecting two such five-qubit arrays by a tunable coupling, assumed to be of the form $\hat{X}^{(4)}\hat{X}^{(7)}$, acting on the fourth and seventh qubit, the Hamiltonian of the composite system is written as
$\hat{H}(t)^{2T_{5}}=\hat{H}^{5Q,\,A}(t)\otimes\mathds{1}+\mathds{1}\otimes\hat{H}^{5Q,\,B}(t)+w_{ent}(t)\hat{X}^{(4)}\hat{X}^{(7)}.$ (4)
@@
Once we have constructed a system that is evolution operator controllable, like the one displayed in Fig. 1(b), we can add more controls without affecting controllability. On the contrary, this can be beneficial, allowing for faster unitary gates or generating the same Lie algebra using commutators of lower depth. As an example, consider the task of implementing an entangling operation on the zeroth and ninth qubit in Fig. 1(b). Intuitively, the fastest option is expected if a tunable coupling between those two qubits exists, whereas the two qubits are at a distance of seven (six static couplings plus the entangling control following the chain 0-1-3-4-7-6-8-9) with the setup of Fig. 1(b). This example illustrates a typical trade-off between connectivity and operation time: The quantum speed limit of gates [38, 39, 40, 41, 42] in large qubit arrays with minimal number of controls and couplings may grossly surpass the decoherence limit of the system.
@@ III Design of large qubit arrays
Next we discuss how to use Theorem 1 to identify possible designs for large qubit arrays with fewer controls and couplings than those currently in use (knowing that the question of operation time will also have to be addressed). To showcase this, we take as basic layout one of the quantum processors from IBM and design a similar architecture using smaller QPUs that have already been shown to be controllable. This will allow us to identify local controls and couplings that can be removed while maintaining evolution operator controllability, i.e., the capability for universal computing.
The start point is a 127-qubit system with similar nearest-neighbor connectivity as shown in Fig. 2, but with local controls on every qubit and tunable two-qubit couplings on every shown connection. This two-dimensional layout is well suited for our approach, as it already has a regular pattern consisting mainly of three-by-five rectangles. The size of this device would make proving controllability using the graph test [20] or the dimensional expressivity test [21] extremely challenging.
The first building block, or ”module”, is a T-shaped five-qubit system. Its Hamiltonian can be written as
@@
Using copies of the systems with $\hat{H}_{5L}(t)$, $\hat{H}_{5L}(t)$ and $\hat{H}_{5L}(t)$, the complete 127-qubit device can be built up. To make the total system controllable, according to Theorem 1, each subsytem must be connected to another one using a tunable coupling. The result is shown in Figure 2, where some of the tunable couplings from the original layout have been “down-graded” to static couplings, effectively reducing the numbers of controls in the system, and some further ones have been removed altogether (the ones depicted with coiled orange lines) without losing controllability. The main achievement, however, is the reduction in the number of local controls to roughly two fifths of the original amount. This is obtained simply by reducing the number of controls in each module by finding controllable four- and five-qubit arrays with any desired controllability test. Comparing exact numbers, IBM’s system is composed of 127 local controls and 144 tunable couplings; the alternative shown in Fig. 2 has only 52 local controls, 101 static couplings, and 25 tunable couplings (marked as necessary couplings). In addition to these 25 couplings, there are a total of 18 extra tunable couplings which can be removed to lower the number of resources or left in place to speed up quantum information transfer. The modified system is a simplified version of the original one with fewer resources.

### v3

@@ I Introduction
The development of quantum hardware is making fast progress over the past few years, bringing us from the NISQ era [preskill2018quantum] closer to fault tolerance [postler2022demonstration] and practical quantum error correction [acharya2025quantum, Bluvstein2025]. At the forefront are architectures based on trapped neutral [Bluvstein2025, Chiu2025] and ionic [postler2022demonstration, ionqbenchmarking, nigmatullin_experimental_2025] atoms as well as superconducting qubits [acharya2025quantum, castelvecchi2023ibm, patra2024efficient, gao2025establishing, menta2025globally, cioni2024conveyor], all of which are expected to scale up further. Important landmarks such as the implementation of fault-tolerant operations [postler2022demonstration] or demonstrations of break-even for quantum error correction [acharya2025quantum, Bluvstein2025] and quantum simulation [manovitz_quantum_2025, nigmatullin_experimental_2025] shift the focus to larger devices. Typically, scalable designs are obtained via a modular approach, linking together multiple smaller quantum processors with limited connectivity rather than developing larger processors [bruzewicz2019trapped, mohseni2024build, huang2024zap, abughanem2025ibm].
@@
One requirement for realizing quantum computing is the ability to perform every unitary operation [divincenzo2000physical]. In control theory, this is termed evolution operator controllability [dAlessandro2021]. Controllability analysis can thus be used as a valuable tool in quantum chip design to ensure that all operations on the device are possible while at the same time simplifying the device [gago2023graph, gago2023determining]. The same concepts that are at the core of controllability analysis can also be employed to the expressivity of parameterized quantum circuits [gago2023determining, allcock2025generating]. Controllability tests can be used to identify the sweet spot balancing universality and the complexity of the system’s architecture [gago2023graph] where complexity refers to qubit-qubit couplings and external qubit controls, i.e., the resources to connect and alter the qubits. Reducing the amount of built-in resources while maintaining the computational capability of the quantum device has not yet been a focus for device design but will be important for further scaling up devices, as is evident just by looking at the number of classical control lines needed to operate today’s most advanced QPUs [postler2022demonstration, acharya2025quantum, Bluvstein2025, Chiu2025, ionqbenchmarking, castelvecchi2023ibm, patra2024efficient, gao2025establishing].
@@
Controllability can be rigorously proven by studying the dynamical Lie algebra of the system, i.e., the algebra generated by a system’s Hamiltonian split into its time-independent drift and all the control operators [dAlessandro2021]. Full rank of the dynamical Lie algebra implies that the system is evolution operator controllable [dAlessandro2021]. However, only in exceptional cases can the dynamical Lie algebra be calculated analytically [Schirmer_2002, Wang2012, PozzoliJPA22, orozco2024quantum, wiersema2024classification, smith2025optimally]. Alternatively, one can determine the rank of the dynamical Lie algebra numerically. The most straightforward way would be to construct a complete basis of the dynamical Lie algebra (which is also a vector space) but this becomes unstable in the required orthogonalization step already for a relatively small number of qubits. Construction of the full algebra can be circumvented by graph methods [BCCS, gago2023graph]. Nevertheless, this approach will eventually fail too, due to the exponential scaling, with the number of qubits, of the size of Hilbert space and thus the dimension of the graph. While the curse of dimensionality can — in principle — be evaded by carrying out the controllability test via a parametric quantum circuit in a hybrid approach combining classical optimization with a quantum device [gago2023determining], it remains yet to be seen how reliably the dimensional expressivity test at the core of this method can be carried out on noisy quantum devices and for larger numbers of qubits. Thus, at present, controllability tests that are generally applicable as well as demonstratedly reliable and scalable are missing. A change in perspective is made possible when proving controllability of a large quantum system from the controllability of its constituent subsystems [zeier2011symmetry, albertini2025quantum, smith2025optimally] .
@@
Here, we leverage this change in perspective for quantum chip design. Instead of attempting to prove controllability for all qubits at once, we show how to build up controllable qubit arrays from smaller, fully controllable subsystems. We can start from building blocks as small as two (or a few) qubits where it is straightforward to prove controllability. Surprisingly, a single controllable entangling link between two qubits in each of the quantum processing subunits is sufficient. Such links are routinely realized in current quantum computing architectures, for example via tunable couplers [niskanen2007quantum, bialczak2011fast, campbell2023modular] for superconducting qubits. While, for a given number of qubits, our approach will not result in the QPU with the smallest possible number of local controls and qubit-qubit couplings, it allows for designs with significantly reduced resources. As an example, we first discuss how two five-qubit arrays that are connected via a tunable coupler result in a controllable ten-qubit array. We then show that this method can be scaled up to generate large devices, thus opening a new route for more resource-efficient device design.
@@ II Controllability of bipartite systems with an entangling control
We briefly recap evolution operator controllability of a quantum system [dAlessandro2021]. Consider a system with Hilbert space dimension $\dim(\mathcal{H})=n$ whose Hamiltonian can be written as
@@
Now, we show how to ensure that controllability is maintained when assembling a larger system by connecting smaller controllable systems. The main idea is rather simple. Assume that we have two, possibly different, qubit arrays “modules”) with given configurations of qubit-qubit couplings and local qubit controls, that are both evolution operator controllable. Then it suffices to connect the two arrays with a control acting on two qubits, one in each module, that is capable of generating entanglement between them, for the resulting multipartite system (composed of the original arrays and the new entangling control) to also be evolution operator controllable. We can keep connecting modules to the original system with one entangling operation per new module and thus design an arbitrarily large system that is evolution operator controllable.
@@
This result is a particular case of more general results on controllability of multipartite systems [zeier2011symmetry, albertini2025quantum]. Our proof of Theorem 1, presented in Appendix A, builds a full basis of the Lie algebra space. This provides information on the depth of the commutators required to implement an operation, in contrast with proofs based on symmetry arguments [zeier2011symmetry] or graph theory [albertini2025quantum]. In more detail, with our new approach, it is possible to determine the number of commutators needed to generate a given element of the Lie algebra, which in turn is related to the mimimal circuit depth required for a given unitary operation.
@@
Considering the simplest example from a gate-oriented perspective, we recover a well-known statement as special case of Theorem 1 — for two qubits, the set of local operations together with any entangling two-qubit gate form a universal set. Here, having access to all local gates implies that the two original systems, the separate qubits, are evolution operator controllable. The entangling gate has the same effect as the entangling control that we use to connect the systems. Note that already in the bipartite case, Theorem 1 is slightly more general: Substitute qudits for the qubits, perhaps of different dimensions $d_{1}=2^{M}$ and $d_{2}=2^{N}$. Since the study is restricted to qubit arrays, the dimension of the qudits is also restricted to powers of 2. The set of local unitary operators for two qubits $SU(2)\otimes SU(2)$ is replaced by $SU(d_{1})\otimes SU(d_{2})$. Every local gate in $SU(d_{1})\otimes SU(d_{2})$ can be implemented given the assumption of controllability on the separate partitions. Theorem 1 then states that given two controllable qudits with dimensions $d_{1}$ and $d_{2}$ and an entangling Hermitian operator $\hat{G}$, the set composed of local operations on the two qudits and the entangling rotation gates $\hat{R}_{\hat{G}}(\phi)$ form a universal set. This universality of the set of local gates plus an entangling gate on qudits of the same size is a lesser known fact in quantum information [muthukrishnan2000multivalued, brylinski2002mathematics, brennen2005criteria].
@@
Once two controllable systems are connected by an entangling control (e.g. a tunable coupler), they can be understood as a controllable system in itself. This implies that we can connect it to a different controllable system using a new entangling control to build a larger yet controllable system. Note that the type of entangling operation between partitions is not relevant to prove controllability. In particular, the statement is true for any two-qubit tunable coupling as long as each of the qubits belongs to a different partition. Such couplings are now routinely used in superconducting qubit architectures [Tan2023tunable, moskalenko2021tunable], including the quantum processors by IBM [miessen2024benchmarking].
@@
Naturally, having only one entangling operation between two partitions will limit the speed at which information (and in particular entanglement) can travel through the system. At the same time, a modular approach will also ease transpilation of quantum algorithms, making the use of partitions that are large enough to run certain subtasks a viable option. This type of architecture is also ideal for using each partition as sub-unit when parallelizing a more complex quantum algorithm. While these use cases would be specifically tailored to the modular design, the universality of the composite system ensures that any kind of quantum circuit can be transpiled into a set of gates suitable for the modular system.
@@
Figure 1: (a): Controllable five-qubit system using just two controls [gago2023graph]. Qubits with (without) local controls are shaded (blank) and static couplings are drawn as straight black lines. (b): Controllable system built up from two copies of the system in (a) with a tunable coupling (blue zig-zag connection).
@@
As an example, consider a T-shaped five-qubit system, inspired by the former five-qubit ibm_quito quantum processor by IBM (cf. https://quantum.cloud.ibm.com/docs/en/guides/retired-qpus), the controllability of which has been proven using a graph method [gago2023graph]. When properly placed, two local controls are sufficient for the five-qubit system with fixed two-qubit couplings in the form of $\hat{H}^{(k,l)}=\hat{\sigma}_{x}^{(k)}\hat{\sigma}_{x}^{(l)}+\hat{\sigma}_{y}^{(k)}\hat{\sigma}_{y}^{(l)}$ to be evolution operator controllable, as illustrated also in Fig. 1(a). Its Hamiltonian is given by
$\hat{H}^{5Q}(t)=$ $-\sum_{j=0}^{4}\frac{\omega_{j}}{2}\hat{\sigma}_{z}^{(j)}+\sum_{\begin{subarray}{c}(k,l)\in\{(0,1),(1,2),\\ (1,3),(3,4)\}\end{subarray}}J_{k,l}\hat{H}^{(k,l)}$
$+u_{1}(t)\hat{\sigma}_{x}^{(1)}+u_{2}(t)\hat{\sigma}_{x}^{(3)}\,$, (3)
@@
Connecting two such five-qubit arrays by a tunable coupling, assumed to be of the form $\hat{\sigma}_{x}^{(4)}\hat{\sigma}_{x}^{(7)}$, acting on the fourth and seventh qubit, the Hamiltonian of the composite system is written as
$\hat{H}(t)^{2T_{5}}=\hat{H}^{5Q,\,A}(t)\otimes\mathds{1}+\mathds{1}\otimes\hat{H}^{5Q,\,B}(t)+w_{ent}(t)\hat{\sigma}_{x}^{(4)}\hat{\sigma}_{x}^{(7)}.$ (4)
@@
Once we have constructed a system that is evolution operator controllable, like the one displayed in Fig. 1(b), we can add more controls without affecting controllability. On the contrary, this can be beneficial, allowing for faster unitary gates or generating the same Lie algebra using commutators of lower depth. As an example, consider the task of implementing an entangling operation on the zeroth and ninth qubit in Fig. 1(b). Intuitively, the fastest option is expected if a tunable coupling between those two qubits exists, whereas the two qubits are at a distance of seven (six static couplings plus the entangling control following the chain 0-1-3-4-7-6-8-9) with the setup of Fig. 1(b). This example illustrates a typical trade-off between connectivity and operation time: The quantum speed limit of gates [DeffnerReview2017, GoerzNPJQI17, GoerzJPB11, GuentherAVSQS2023, BasilewitschPRR2024] in large qubit arrays with minimal number of controls and couplings may grossly surpass the decoherence limit of the system.
@@ III Design of large qubit arrays
Next we discuss how to use Theorem 1 to identify possible designs for large qubit arrays with fewer controls and couplings than those currently in use (knowing that the question of operation time will also have to be addressed). To showcase this, we take as basic layout one of the quantum processors from IBM, composed of 127 qubits. Several quantum algorithms have been implemented on this platform, using at least a hundred qubits. These include quantum simulations with 2D tranverse-field Ising models [kim2023evidence] and preparing the ground state of the lattice Schwinger model [farrell2024scalable], as two examples showcasing the benefit of controllable qubit arrays of increasing size. The goal is then to design a similar architecture using smaller QPUs that have already been shown to be controllable. This will allow us to identify local controls and couplings that can be removed while maintaining evolution operator controllability, i.e., the capability for universal computing.
The start point is a 127-qubit system with similar nearest-neighbor connectivity as shown in Fig. 2, but with local controls on every qubit and tunable two-qubit couplings on every shown connection. This two-dimensional layout is well suited for our approach, as it already has a regular pattern consisting mainly of three-by-five rectangles. The size of this device would make proving controllability using the graph test [gago2023graph] or the dimensional expressivity test [gago2023determining] extremely challenging.
The first building block, or “module”, is a T-shaped five-qubit system. Its Hamiltonian can be written as
@@
Using copies of the systems with $\hat{H}_{5T}(t)$, $\hat{H}_{5L}(t)$ and $\hat{H}_{4L}(t)$, the complete 127-qubit device can be built up. To make the total system controllable, according to Theorem 1, each subsytem must be connected to another one using a tunable coupling. The result is shown in Figure 2, where some of the tunable couplings from the original layout have been “down-graded” to static couplings, effectively reducing the numbers of controls in the system, and some further ones have been removed altogether (the ones depicted with coiled orange lines) without losing controllability. The main achievement, however, is the reduction in the number of local controls to roughly two fifths of the original amount. This is obtained simply by reducing the number of controls in each module by finding controllable four- and five-qubit arrays with any desired controllability test. Comparing exact numbers, IBM’s system is composed of 127 local controls and 144 tunable couplings; the alternative shown in Fig. 2 has only 52 local controls, 101 static couplings, and 25 tunable couplings (marked as necessary couplings). In addition to these 25 couplings, there are a total of 18 extra tunable couplings which can be removed to lower the number of resources or left in place to speed up quantum information transfer. The modified system is a simplified version of the original one with fewer resources.